# RealCam-I2V: Real-World Image-to-Video Generation with Interactive Complex Camera Control

### v1

@@ Abstract
Recent advancements in camera-trajectory-guided image-to-video generation offer higher precision and better support for complex camera control compared to text-based approaches. However, they also introduce significant usability challenges, as users often struggle to provide precise camera parameters when working with arbitrary real-world images without knowledge of their depth nor scene scale. To address these real-world application issues, we propose RealCam-I2V, a novel diffusion-based video generation framework that integrates monocular metric depth estimation to establish 3D scene reconstruction in a preprocessing step. During training, the reconstructed 3D scene enables scaling camera parameters from relative to absolute values, ensuring compatibility and scale consistency across diverse real-world images. In inference, RealCam-I2V offers an intuitive interface where users can precisely draw camera trajectories by dragging within the 3D scene. To further enhance precise camera control and scene consistency, we propose scene-constrained noise shaping, which shapes high-level noise and also allows the framework to maintain dynamic, coherent video generation in lower noise stages. RealCam-I2V achieves significant improvements in controllability and video quality on the RealEstate10K and out-of-domain images. We further enables applications like camera-controlled looping video generation and generative frame interpolation. We will release our absolute-scale annotation, codes, and all checkpoints. Please see dynamic results in https://zgctroy.github.io/RealCam-I2V.
@@
Figure 2: Comparision between text, relative trajectory, and absolute trajectory based camera-controlled image-to-video generation methods on aspects of camera control precision and usability.
@@ 1 Introduction
Recent advancements in image-to-video generation [14, 18, 7, 8, 4, 66] have significantly improved controllability over synthesized videos. However, challenges remain in achieving realistic, controllable camera movement within complex real-world scenes. Text-based camera-control methods [14, 2, 28, 22, 24, 50], like traditional diffusion-based video generation, are intuitive and straightforward but lack precision in explicit control over camera parameters, such as angle, scale, and movement direction. This limitation has spurred the development of camera-trajectory-guided approaches, which attempt to address these issues by offering finer control over camera movement.
@@
Current camera-trajectory-guided methods typically rely on relative camera trajectories, as seen in models like MotionCtrl [53], CameraCtrl [16], CamCo [62], and CamI2V [77]. While these approaches provide more control than text-based models, they are fundamentally limited by their reliance on relative scale trajectories. Training on relative scales results in inconsistencies when applied to real-world scenes, where absolute scale is crucial for realistic depth perception. Additionally, without access to depth information, users find it challenging to draw precise trajectories, making these methods difficult to use effectively.
@@
To overcome these limitations, we propose RealCam-I2V, a video generation framework that integrates monocular depth estimation as a preprocessing step to construct a robust, absolute-scale 3D scene. Our approach leverages the Depth Anything v2 [64] model to predict the metric depth of a user-provided reference image, reprojecting its pixels back into camera space to create a stable 3D representation. This 3D scene serves as the foundation for camera control, providing a consistent and absolute scale that is critical for real-world applications.
@@
In the training stage, we align the reconstructed 3D scene of the reference image with the point cloud of each video sample, reconstructed using COLMAP [40], a structure-from-motion (SfM) method. This alignment allows us to rescale COLMAP-annotated camera parameters to the Depth Anything metric, providing an absolute, stable, and robust scale across training data. By aligning relative-scale camera parameters to absolute scales, we can condition the video generation model on accurately scaled camera trajectories, achieving greater control and scene consistency across diverse real-world images.
@@
During inference, RealCam-I2V provides an interactive interface where users can intuitively design camera trajectories by drawing within the reconstructed 3D scene of the reference image. This interface renders preview videos of the trajectory in a static scene, offering users real-time feedback and greater control over camera movement. This interactive feature enhances usability, allowing precise trajectory control even for users without specialized knowledge of scene depth. To further improve video quality and control precision, we introduce scene-constrained noise initialization as a mechanism to shape the generation process in its high-noise stages. By using the preview video of the static 3D scene, RealCam-I2V injects scene-visible regions with controlled noise, guiding the video diffusion model’s early generation stages. This high-noise feature constrains the initial layout and camera dynamics, providing a strong foundation for the remaining denoising stages. As denoising progresses, the condition-based approach, trained on absolute-scale camera trajectories, preserves global layout and completes the dynamic scene in previously unseen regions. This approach maintains the video diffusion model’s capacity for dynamic content generation while ensuring accurate, coherent camera control.
@@
Our experimental results show that RealCam-I2V achieves significant performance gains in video quality and controllability. When relative scales are aligned to absolute scales, models such as MotionCtrl, CameraCtrl, and CamI2V see substantial improvements in video quality. Furthermore, with the introduction of scene-constrained noise initialization, RealCam-I2V surpasses state-of-the-art performance benchmarks, particularly on datasets like RealEstate10K [78] and out-of-domain images. These results demonstrate the effectiveness of our approach in both controlled and diverse real-world settings. In summary, our contributions are as follows:
@@
  • With reconstructed 3D scene, we enable absolute-scale training and provide an interactive interface during inference to easily design camera trajectories with preview feedback, along with proposed scene-constrained noise shaping to significantly enhance scene consistency and camera controllability.
@@ 2 Related Works
Diffusion-based Video Generation. The advancement of diffusion models [39, 38, 75] has led to significant progress in video generation. Due to the scarcity of high-quality video-text datasets [2, 3], researchers have adapted existing text-to-image (T2I) models to facilitate text-to-video (T2V) generation. Notable examples include AnimateDiff [14], Align your Latents [3], PYoCo [11], and Emu Video [12]. Further advancements, such as LVDM [18], VideoCrafter [7, 8], ModelScope [47], LAVIE [51], and VideoFactory [49], have refined these approaches by fine-tuning both spatial and temporal blocks, leveraging T2I models for initialization to improve video quality. Recently, Sora [4] and CogVideoX [66] enhance video generation by introducing Transformer-based diffusion backbones [36, 30, 69] and utilizing 3D-VAE, unlocking the potential for realistic world simulators. Additionally, SVD [2], SEINE [9], PixelDance [70] and PIA [74] have made significant strides in image-to-video generation, achieving notable improvements in quality and flexibility. Further, I2VGen-XL [73], DynamicCrafter [59], and Moonshot [71] incorporate additional cross-attention layers to strengthen conditional signals during generation.
Controllable Generation. Controllable generation has become a central focus in both image [72, 25, 33, 76, 37, 67, 56, 42, 58] and video [13, 71, 15, 26] generation, enabling users to direct the output through various types of control. A wide range of controllable inputs has been explored, including text descriptions, pose [31, 48, 21, 63], audio [43, 44, 17], identity representations [5, 52, 57], trajectory [68, 6, 29, 55, 34].
@@
Trajectory-based Camera Control. MotionCtrl [53], CameraCtrl [16], and Direct-a-Video [65] use camera pose as input to enhance control, while CVD [27] extends CameraCtrl for multi-view generation, though still limited by motion complexity. To improve geometric consistency, Pose-guided diffusion [45], CamCo [61], and CamI2V [77] apply epipolar constraints for consistent viewpoints. VD3D [1] introduces a ControlNet[72]-like conditioning mechanism with spatiotemporal camera embeddings, enabling more precise control. CamTrol [19] offers a training-free approach that renders static point clouds into multi-view frames for video generation. Cavia [60] introduces view-integrated attention mechanisms to improve viewpoint and temporal consistency, while I2VControl-Camera [10] refines camera movement by employing point trajectories in the camera coordinate system. Despite these advancements, challenges in maintaining camera control and scene-scale consistency remain, which our method seeks to address. It is noted that 4Dim [54] introduces absolute scale but in 4D novel view synthesis (NVS) of scenes.
Figure 3: Pipeline. In training, we align camera parameters in RealEstate10K from relative scale to absolute scale. In inference, we use metric depth estimation method to construct a 3d point cloud for users to interactively drawing camera traces.
Figure 4: 3D point cloud reconstructed from metric depth estimation (RGB) is robust and unified, whereas the SFM-based reconstruction by methods like COLMAP (Yellow) used in RealEstate10K annotations is in relative scale and may vary across images. Aligning these two 3D scenes enables the transformation from relative to absolute scale (real-world scale).
@@ 3 Method
3.1 Metric Depth Estimation for 3D Reconstrcution
To obtain a depth map from a given input image, we use a metric depth predictor $f_{\text{depth}}$, which takes the RGB image $I$ as input and outputs the corresponding depth map $D(u,v)$. The prediction process is formulated as:
$D(u,v)=f_{\text{depth}}(I),$
$K=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix},$
where $f_{x}$ and $f_{y}$ are the focal lengths along the $x$ and $y$ axes, $(c_{x},c_{y})$ is the principal point of the camera. Given a depth map $D(u,v)$, the projected 3D coordinates in the camera coordinate system, $\mathbf{p}_{c}=(x_{c},y_{c},z_{c})^{T}$, are computed as:
$\mathbf{p}_{c}=D(u,v)\cdot K^{-1}\cdot\begin{bmatrix}u\\ v\\ 1\end{bmatrix}.$
3.2 Absolute-Scale Training
Camera-controlled Image-to-Video Model. Instead of directly modeling the video $x$, the latent representation $z=\mathcal{E}(x)$ is used for training. The diffusion model $\epsilon_{\theta}$ learns to estimate the noise $\epsilon$ added at each timestep $t$, conditioned on both a text prompt $c_{\text{txt}}$, a reference image $c_{\text{img}}$, and camera condition $c_{\text{cam}}$, with $t\in\mathcal{U}(0,1)$. The training objective simplifies to a reconstruction loss defined as:
@@ 3.2 Absolute-Scale Training
Aligning from Relative Scale to Absolute Scale. To convert camera extrinsics from world-to-camera to an absolute-scale camera-to-world representation, we defines that the world-to-camera extrinsics matrix $F_{\text{w2c}}\in\mathbb{R}^{4\times 4}$ is inverted to obtain the corresponding camera-to-world matrix:
$F_{\text{c2w}}=F_{\text{w2c}}^{-1}.$
To express the transformations relative to the first frame, each $F_{\text{c2w}}$ is left-multiplied by the camera-to-world matrix of the inverse of first frame $F_{\text{c2w, 1}}$:
@@
Here, $c_{\text{cam}}\in\mathbb{R}^{F\times 4\times 4}$ represents the camera-to-world transformations aligned relative to the first frame. However, the translation component of $c_{\text{cam}}$ remains in a relative scene scale. To convert the relative translation to an absolute scale, we align the metric 3D point cloud reconstructed by Depth Anything with the 3D point cloud reconstructed by COLMAP (Structure-from-Motion), as shown in Fig. 4. The alignment process yields a scale factor $a$ and is applied to the translation component of $c_{\text{cam}}$, resulting in an absolute-scale camera-to-world transformation:
$c_{\text{cam}}^{\text{abs}}=\begin{bmatrix}R&a\cdot T\\ 0&1\end{bmatrix},$
@@
where $R$ is the rotation matrix, $T$ is the relative translation vector. The resulting $c_{\text{cam}}^{\text{abs}}\in\mathbb{R}^{F\times 4\times 4}$ represents the camera-to-world transformations with absolute scene scale, enabling robust and accurate real-world applications.
3.3 Scene-Constrained Noise Shaping
Inspired by SDEdit [32] and DDIM [41] inversion, noised features $z_{t}$ can be used for shaping the layout, camera control of the entire image, especially at timestep with high-level noise. We propose scene-constrained noise shaping, which utilizes preview videos generated along user-defined trajectories in the interactive 3D scene. Each frame of the preview video is treated as a reference frame and provided to the generation process during the high-noise stage. The reference frame’s pixels are overlaid onto the model-predicted $z_{0}$ to achieve the shaping effect.
Next, we detail the process for selecting the pixels to be referenced. As illustrated in Fig. 5, the primary criterion is that a pixel must be visible under the current camera viewpoint in the preview video. To mitigate issues such as holes caused by inaccurate depth predictions, we apply an additional filtering rule: if a visible pixel’s $k\times k$ neighborhood contains any invisible pixels, it is considered to lie on an object’s edge and potentially affected by depth prediction errors. Such pixels are excluded from selection. Finally, we define the noise shaping process with the following formula:
$z_{\text{predict}}=\text{mask}\cdot z_{\text{preview}}+(1-\text{mask})\cdot z_{\text{predict}},$
where the mask identifies the selected reference pixels, $z_{\text{preview}}$ represents the latent features from the preview video, and $z_{\text{predict}}$ is the model-predicted latent representation.
@@ 3.3 Scene-Constrained Noise Shaping
Figure 5: Pixels selected for scene-constrained noise shaping, where they are pasted onto the clean part (predicted $z_{0}$) of a noised latent $z_{t}$, typically at high noise levels $1.00.9<t<$ 1.0 is enough for camera control and maintain dynamics.
@@
Figure 6: Concatenation for different tasks, including basic mode, interpolation mode, and continuation mode.
3.4 Interploatation, Loop and Continueation

### v2

@@ Abstract
Recent advancements in camera-trajectory-guided image-to-video generation offer higher precision and better support for complex camera control compared to text-based approaches. However, they also introduce significant usability challenges, as users often struggle to provide precise camera parameters when working with arbitrary real-world images without knowledge of their depth nor scene scale. To address these real-world application issues, we propose RealCam-I2V, a novel diffusion-based video generation framework that integrates monocular metric depth estimation to establish 3D scene reconstruction in a preprocessing step. During training, the reconstructed 3D scene enables scaling camera parameters from relative to metric scales, ensuring compatibility and scale consistency across diverse real-world images. In inference, RealCam-I2V offers an intuitive interface where users can precisely draw camera trajectories by dragging within the 3D scene. To further enhance precise camera control and scene consistency, we propose scene-constrained noise shaping, which shapes high-level noise and also allows the framework to maintain dynamic and coherent video generation in lower noise stages. RealCam-I2V achieves significant improvements in controllability and video quality on the RealEstate10K and out-of-domain images. We further enables applications like camera-controlled looping video generation and generative frame interpolation. Project page: zgctroy.github.io/RealCam-I2V.
@@
11footnotetext: Equal Contribution
@@ 1 Introduction
Recent advancements in image-to-video generation [20, 24, 9, 10, 83] have significantly improved controllability over synthesized videos. However, challenges remain in achieving realistic, controllable camera movement within complex real-world scenes. Text-based camera-control methods [20, 4, 38, 29, 31, 66], like traditional diffusion-based video generation, are intuitive and straightforward but lack precision in explicit control over camera parameters, such as angle, scale, and movement direction. This limitation has spurred the development of camera-trajectory-guided approaches, which attempt to address these issues by offering finer control over camera movement.
@@
Current camera-trajectory-guided methods typically rely on relative camera trajectories, as seen in models like MotionCtrl [69], CameraCtrl [22], CamCo [79], and CamI2V [97]. While these approaches provide more control than text-based models, they are fundamentally limited by their reliance on relative scale trajectories. Training on relative scales results in inconsistencies when applied to real-world scenes, where metric scale is crucial for realistic depth perception. Additionally, without access to depth information, users find it challenging to draw precise trajectories, making these methods difficult to use effectively.
@@
To overcome these limitations, we propose RealCam-I2V, a video generation framework that integrates monocular depth estimation as a preprocessing step to construct a robust, metric-scale 3D scene. Our approach leverages the Depth Anything v2 [81] model (metric version) to predict the metric depth of a user-provided reference image, reprojecting its pixels back into camera space to create a stable 3D representation. This 3D scene serves as the foundation for camera control, providing a consistent and metric scale that is critical for real-world applications.
@@
In the training stage, we align the reconstructed 3D scene of the reference image with the point cloud of each video sample, reconstructed using COLMAP [54], a structure-from-motion (SfM) method. This alignment allows us to rescale COLMAP-annotated camera parameters to the Depth Anything metric, providing an metric, stable, and robust scale across training data. By aligning relative-scale camera parameters to metric scales, we can condition the video generation model on accurately scaled camera trajectories, achieving greater control and scene consistency across diverse real-world images.
@@
During inference, RealCam-I2V provides an interactive interface where users can intuitively design camera trajectories by drawing within the reconstructed 3D scene of the reference image. This interface renders preview videos of the trajectory in a static scene, offering users real-time feedback and greater control over camera movement. This interactive feature enhances usability, allowing precise trajectory control even for users without specialized knowledge of scene depth. To further improve video quality and control precision, we introduce scene-constrained noise initialization as a mechanism to shape the camera movement in its high-noise stages. By using the preview video of the static 3D scene, RealCam-I2V injects scene-visible regions with controlled noise, guiding the video diffusion model’s early generation stages. This high-noise feature constrains the initial layout and camera dynamics, providing a strong foundation for the remaining denoising stages. As denoising progresses, the condition-based approach, trained on metric-scale camera trajectories, preserves global layout and completes the dynamic scene in previously unseen regions. This approach maintains the video diffusion model’s capacity for dynamic content generation while ensuring accurate, coherent camera control.
@@
Our experimental results show that RealCam-I2V achieves significant performance gains in video quality and controllability. When relative scales are aligned to metric scales, models such as MotionCtrl, CameraCtrl, and CamI2V see substantial improvements in video quality. Furthermore, with the introduction of scene-constrained noise initialization, RealCam-I2V surpasses state-of-the-art performance benchmarks, particularly on datasets like RealEstate10K [101] and out-of-domain images. These results demonstrate the effectiveness of our approach in both controlled and diverse real-world settings. In summary, our contributions are as follows:
@@
  • With reconstructed 3D scene, we enable metric-scale training and provide an interactive interface during inference to easily design camera trajectories with preview feedback, along with proposed scene-constrained noise shaping to significantly enhance scene consistency and camera controllability.
@@
Figure 2: RealCam-I2V pipeline. For training, we align camera parameters from relative scale to metric scale. For inference, we use metric depth estimation to construct the point cloud for users to interactively draw the camera trajectory. Due to the metric scale alignment, the user-given camera trajectory in the 3D scene shares the same scene scale as those in real world.
@@ 2 Related Works
Diffusion-based Video Generation. The advancements in diffusion models [53, 52, 95] have led to significant progress in video generation, gradually forming two major categories of foundational tasks: text-to-video (T2V) and image-to-video (I2V). Due to the lack of high-quality video-text datasets [4, 5], previous outstanding works such as AnimateDiff [20], Align Your Latents [5], PYoCo [17], Emu Video [18], LVDM [24], VideoCrafter [9, 10], ModelScope [62], LAVIE [67], and VideoFactory [65] have facilitated text-to-video (T2V) generation by incorporating Motion Blocks into existing text-to-image [51, 53, 86] models. Building on these T2V efforts, several works introduced images as control signals, leading to the development of a series of I2V models. Representative examples include SVD [4], SEINE [13], PixelDance [88], PIA [93], I2VGen-XL [92], DynamicCrafter [76], and Moonshot [89]. With the improvement in video data [12, 47, 63, 34] and the rapid development of the DiT architecture [49, 42, 87], recent works such as CogVideoX [83], Sora [6], HunyuanVideo [35], GoKu [11], Open-Sora [99], Open-Sora-Plan [40], Vidu [3], Lumina-T2X [16], Vchitect-2.0 [14], RepVideo [55] and Step-Video-T2V[41] have significantly enhanced the quality of video generation. These advancements demonstrate the potential of Video Diffusion Models as realistic world simulators.
Controllable Generation. Controllable generation has become a central focus in both image [91, 32, 45, 96, 50, 84, 72, 57, 75] and video [19, 89, 21, 33] generation, enabling users to direct the output through various types of control. A wide range of controllable inputs has been explored, including text descriptions, pose [43, 64, 27, 80], audio [58, 59, 23], identity representations [7, 68, 74, 73, 94, 37], trajectory [85, 8, 39, 71, 46].
@@
Trajectory-based Camera Control. MotionCtrl [69], CameraCtrl [22], and Direct-a-Video [82] use camera pose as input to enhance control, while CVD [36] extends CameraCtrl for multi-view generation, though still limited by motion complexity. To improve geometric consistency, Pose-guided diffusion [60], CamCo [78], and CamI2V [97] apply epipolar constraints for consistent viewpoints. VD3D [2] introduces a ControlNet[91]-like conditioning mechanism with spatiotemporal camera embeddings, enabling more precise control. CamTrol [25] offers a training-free approach that renders static point clouds into multi-view frames for video generation. Cavia [77] introduces view-integrated attention mechanisms to improve viewpoint and temporal consistency, while I2VControl-Camera [15] refines camera movement by employing point trajectories in the camera coordinate system. Recently, 4DiM [70] and AC3D [1] also leverage monocular metric depth estimator to tackling the issues of scale inconsistency. Despite these advancements, challenges in maintaining camera control and scene-scale consistency remain, which our method seeks to address.
@@ 3 Method
Our overall pipeline is shown in Fig. 2. For training, we align camera parameters from relative scale to metric scale. For inference, we use metric depth estimation to construct the point cloud for users to interactively draw the camera trajectory. Due to the metric scale alignment, the user-given camera trajectory in the 3D scene shares the same scene scale as those in real world.
3.1 Camera-controlled Image-to-Video Model
Instead of directly modeling the video $x$, the latent representation $z=\mathcal{E}(x)$ is used for training. The diffusion model $\epsilon_{\theta}$ learns to estimate the noise $\epsilon$ added at each timestep $t$, conditioned on both a text prompt $c_{\text{txt}}$, a reference image $c_{\text{img}}$, and camera condition $c_{\text{cam}}$, with $t\in\mathcal{U}(0,1)$. The training objective simplifies to a reconstruction loss defined as:
@@ 3.1 Camera-controlled Image-to-Video Model
Figure 3: Scene scale mismatch. Point clouds reconstructed from metric depth estimation (RGB) are robust and unified, whereas SfM reconstructions (yellow) are relative-scale that may vary across frames. Our alignment enables relative-to-metric conversion of scene scale for real-world applications.
Figure 4: Camera trajectory ambiguity. The relative scene-scale measurement fundamentally hinders models from learning physically consistent camera motion.
3.2 Metric Scene-scale Alignment
Metric Depth Estimation. To obtain a depth map from a given input image, we use a metric depth predictor $f_{\text{depth}}$, which takes the RGB image $I$ as input and outputs the corresponding depth map $D(u,v)$. The prediction process is formulated as:
$D(u,v)=f_{\text{depth}}(I),$
$K=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix},$
where $f_{x}$ and $f_{y}$ are the focal lengths along the $x$ and $y$ axes, $(c_{x},c_{y})$ is the principal point of the camera.
Metric Scene-Scale Alignment. To convert camera extrinsics from world-to-camera to an metric-scale camera-to-world representation, we defines that the world-to-camera extrinsics matrix $F_{\text{w2c}}\in\mathbb{R}^{4\times 4}$ is inverted to obtain the corresponding camera-to-world matrix $F_{\text{c2w}}=F_{\text{w2c}}^{-1}.$ To express the transformations relative to the first frame, each $F_{\text{c2w}}$ is left-multiplied by the camera-to-world matrix of the inverse of first frame $F_{\text{c2w, 1}}$:
@@ 3.2 Metric Scene-scale Alignment
Here, $c_{\text{cam}}\in\mathbb{R}^{F\times 4\times 4}$ represents the camera-to-world transformations aligned relative to the first frame. However, the translation component of $c_{\text{cam}}$ remains in a relative scene scale. To convert the relative translation to an metric scale, we align the metric 3D point cloud reconstructed by Depth Anything (metric version) with the 3D point cloud reconstructed by COLMAP (Structure-from-Motion), as shown in Fig. 3 and Fig. 4. The alignment process yields a scale factor $\alpha$ and is applied to the translation component of $c_{\text{cam}}$, resulting in an metric-scale camera-to-world transformation:
$c_{\text{cam}}^{\text{metric}}=\begin{bmatrix}R&\alpha\cdot T\\ 0&1\end{bmatrix},$
@@
where $R$ is the rotation matrix, $T$ is the relative translation vector. The resulting $c_{\text{cam}}^{\text{metric}}\in\mathbb{R}^{F\times 4\times 4}$ represents the camera-to-world transformations with metric scene scale, enabling robust and accurate real-world applications.
Figure 5: One-round generation versus multi-round generation. Our framework decouples camera adjustment via interactive 3D scenes, enabling fast feedback before slow generation.
3.3 One-round Generation in Inference
Users are allowed to draw camera trajectory in the constructed 3D scene and can easily preview the rendering video of expected camera trajectory, free from the costly generation via video diffusion models, as shown in Fig. 5.
3D Point Projection For Interaction. Given a depth map $D(u,v)$, the projected 3D coordinates in the camera coordinate system, $\mathbf{p}_{c}=(x_{c},y_{c},z_{c})^{T}$, are computed as:
$\mathbf{p}_{c}={\bf D}(u,v)\cdot K^{-1}\cdot\begin{bmatrix}u\\ v\\ 1\end{bmatrix}.$
@@ 3.3 One-round Generation in Inference
Figure 6: Pixel selection for noise shaping. Selected pixels are pasted onto the clean part (predicted $z_{0}$) of noised latent $z_{t}$. Empirically, applying noise shaping at high noise levels $t>0.9$ achieves the trade-off of camera control and dynamics.
Preview Video Rendering for Camera Motion. With constructed 3D point cloud, we can render a preview video of camera traces. As shown in Fig. 7, the preview video functions as a reference video for validation.
Scene-constrained Noise Shaping. Inspired by previous works SDEdit [44] and DDIM inversion [56], noised features $z_{t}$ can be used for shaping the layout, camera control of the entire image, especially at timestep with high-level noise. We propose scene-constrained noise shaping, which utilizes preview videos generated along user-defined trajectories in the interactive 3D scene. Each frame of the preview video is treated as a reference frame and provided to the generation process during the high-noise stage. The reference frame’s pixels are overlaid onto the model-predicted $z_{0}$ to achieve the shaping effect.
Next, we detail the process for selecting the pixels to be referenced. As illustrated in Fig. 6, the primary criterion is that a pixel must be visible under the current camera viewpoint in the preview video. To mitigate issues such as holes caused by inaccurate depth predictions, we apply an additional filtering rule: if a visible pixel’s $k\times k$ neighborhood contains any invisible pixels, it is considered to lie on an object’s edge and potentially affected by depth prediction errors. Such pixels are excluded from selection. Finally, we define the noise shaping process as the formula:
$z_{t}=m\cdot(\alpha_{t}z_{\text{preview}}+\sigma_{t}\epsilon)+(1-m)\cdot z_{t},$
where $m$ identifies the selected reference pixels, $z_{\text{preview}}$ represents the clean latent features from the preview video. It should be noted that resampling $\epsilon$ in each timestep is important because the motivation of noise shaping is to utilize the clean feature of preview feature to shape the low-frequency information of the generation process while the random sampled $\epsilon$ may cover some of the valid information in each timestep if it is sampled to be a large value.
Figure 7: Preview video rendering for user-expected camera trajectory. It also serves as the reference video in our proposed scene-constrained noise shaping.
@@
Figure 8: Concatenation strategies for different tasks. Basic mode, interpolation mode, and continuation mode can be supported with only minor modification.
Relative Scale Metric Scale Relative Scale Metric Scale VideoGPT StyleGAN
  DynamiCrafter [76] 3.3415 9.8024 14.135 11.625 15.726 106.02 92.196
MotionCtrl [69] 1.0527 2.2860 6.8182 2.9312 7.2272 70.292 60.845
CameraCtrl [22] 0.7373 1.7619 5.5090 2.1644 5.7648 69.202 58.900
CamI2V [97] 0.4120 1.3409 3.2934 1.5291 3.4230 62.439 53.361
RealCam-I2V (Ours) 0.3884 1.2943 2.2317 1.4628 2.3609 53.718 45.460
w.r.t. CamI2V +5.73% +3.48% +32.24% +4.33% +31.03% +13.97% +14.81%
Table 1: Quantitative comparison with SOTA methods. Our approach excels all baselines on both relative and metric results, while coherently improve visual quality of generated videos. We observe over 30% improvement on metric scale results and over 10% improvement on FVD. * denotes our reproduced results on DynamiCrafter [76]. Best and second best results are highlighted respectively.
3.4 Interpolation, Loop and Continuation